meera mother is four times as old as meera. after 5 years her mother will be three times as old as she will be then what are their present age.
Found 2 solutions by checkley71, Adam:
Answer by checkley71(8403)

MOTHER=4 MEERA
M+5=3(C+5)
4C+5=3C+15
4C-3C=15-5
C=10 AGE OF MEERA NOW.
PROOF
MOTHER=4*10=40 AGE OF THE MOTHER NOW.
40+5=3(10+5)
45=3*15
45=45

Answer by Adam(64)

We denote mother's age by x. And Meera's age by y. We can write 4*y = x or x/4=y the second equation might be (x+5)/3=y+5 or 3*(y+5)=x+5. Let's choose second option and create a system of 2 linear equations with 2 variables.
4*y = x
3*(y+5)=x+5
We can substitute x in terms of y into second equation, so we get the equation with only variable.
3*(y+5)=4*y+5 multiply
3*y+15=4*y+5 subtract 3*y
15=4*y-3*y+5
15=y+5 /-5
10=y
Now we know that Meera is presently 10 years old.
We can substitute our result into any equation of the original system.
4*10 = x
40= x
Mother's present age is 40.